# Fovea Stacking: Imaging with Dynamic Localized Aberration Correction

### v1

@@ Abstract.
The desire for cameras with smaller form factors has recently lead to a push for exploring computational imaging systems with reduced optical complexity such as a smaller number of lens elements. Unfortunately such simplified optical systems usually suffer from severe aberrations, especially in off-axis regions, which can be difficult to correct purely in software.
In this paper we introduce Fovea Stacking, a new type of imaging system that utilizes emerging dynamic optical components called deformable phase plates (DPPs) for localized aberration correction anywhere on the image sensor. By optimizing DPP deformations through a differentiable optical model, off-axis aberrations are corrected locally, producing a foveated image with enhanced sharpness at the fixation point - analogous to the eye’s fovea. Stacking multiple such foveated images, each with a different fixation point, yields a composite image free from aberrations. To efficiently cover the entire field of view, we propose joint optimization of DPP deformations under imaging budget constraints. Due to the DPP device’s non-linear behavior, we introduce a neural network-based control model for improved alignment between simulation-hardware performance.
@@
ccs: Computing methodologies Computational photography
Figure 1. Fovea Stacking is new computational camera paradigm aimed at simplifying (and ultimately miniaturizing) camera optics. a) the optical system consists of an achromatic doublet lens, which is highly aberrated, and a refractive deformable phase plate (DPP), which can be used to correct the aberrations for a local region of interest (the fovea). This simple optical system does not allow for the simultaneous correction of all aberrations everywhere in the image, however, the fovea can be dynamically moved anywhere in the image by controlling the DPP. b) a foveated image captured with our system, where the fovea is placed in the blue box near the top. c) magnified crop regions for different foveated images. The left column corresponds the image shown in b); the other columns have the fovea moved to the other respective boxes. d) by stacking up the sharpest regions from different foveated images (i.e. a fovea stack), an un-aberrated all-in-focus image can be recovered. Other applications such as object tracking and depth-dependent focusing are discussed in the text.
[…]
presenting the general idea
@@ 1. Introduction
In this work we use the ability of the DPP to dynamically and locally shape an optical wavefront in a compact optical system to demonstrate a new type of foveated imaging system (see Fig. 1). The optical design consists of a combination of a highly aberrated achromatic double lens and a DPP that can be used for aberration correction. This optical system lacks the complexity to simultaneously correct for all aberrations everywhere in the image like a classical camera lens, however, it has the ability to correct for the aberrations in a localized region of interest (the fovea) and produce an image with excellent fidelity in that region. Moreover, through dynamic control of the DPP, this fovea region can be placed anywhere in the image plane without mechanical motion. A stack of such foveated images, a.k.a. a Fovea Stack, can be used to reconstruct an unabberrated, high fidelity image (c.f. Fig. 1d).
@@ 2.1. Wavefront Modulators
Recently, a novel optofluidic refractive wavefront modulator called deformation phase plate (DPP) was proposed by (Banerjee et al., 2018). It consists of an optically transparent liquid-filled cavity sealed by an elastic polymer membrane. The membrane’s deformation is controlled via electrostatic actuation using 63 transparent electrodes on the back plane of the cavity, enabling distortion correction up to seventh radial Zernike order (Rajaeipour et al., 2021). In this work, we adopt the DPP as a refractive wavefront modulator to create a compact imaging system. However, it is observed that traditional linear control strategies are inaccurate for large wavefront deviations. To address this, we propose a neural network-based control strategy to improve accuracy.
Compared to adaptive optics (AO) systems (Hampson et al., 2021), which typically correct for aberrations introduced by atmospheric turbulence or other external media before light reaches the imaging system, our method focuses on correcting distortions caused by the inherent imperfections of simple, compact imaging optics.
@@ 2.2. Focus Stacking
Focus stacking extends depth of field by capturing multiple images at varying focus distances. From this image stack, a sharp all-in-focus (AIF) image can be reconstructed. Similarly, fovea stacking aims to reconstruct a sharp, aberration- corrected image from a stack of regionally sharp images. Focus stacking can also be used correct for field curvature, where the focal region is not planar but a curved manifold. It can therefore serve as a baseline for our work.
Traditional multi-focus image fusion methods reconstruct all-in-focus images by identifying and compositing the sharpest regions from multiple input images  (Huang and Jing, 2007; Nayar and Nakagawa, 1994). This fusion can occur in the spatial domain or in transformed domains, such as Laplacian pyramids  (Wang and Chang, 2011; Sun et al., 2018) and discrete wavelet transforms  (Pradnya and Sachin, 2013). For comprehensive reviews of traditional and deep learning-based methods, see  (Zafar et al., 2020; Zhang, 2021). Alternatively, some approaches jointly reconstruct both a depth map and an all-in-focus image - for example, Wang et al. (Wang et al., 2021) use 3D convolution to learn attention weights from the focus stack, which are then used to produce both the AIF image and the corresponding depth map.
Traditional focus adjustment often involves mechanical movement of the sensor or lens elements, which can induce vibrations during large-amplitude, high-frequency motion  (Kuthirummal et al., 2010). As an alternative, focus-tunable liquid lens (FTLs) have been adopted for focus-sweeping applications  (Miau et al., 2013), but they are limited to tuning optical power without ability to correct higher-order aberrations. Although the adjustable range of the DPP is limited, it is capable to adjust the focus distance by modulating the ”defocus” term in the Zernike polynomial. Our method, therefore, supports stacking not only in the x-y direction but also along the depth direction, enabling aberration-corrected imaging for extended depth-of-field applications.
@@ 2.3. Foveated Imaging
Foveated imaging, inspired by human vision, concentrates high resolution where it is most needed - similar to the function of the fovea. Many different approaches have been proposed to achieve spatially varying resolution through specialized sensor design or image fusion techniques. Retina-like sensors, for example, arrange photoreceptors in a spatially varying log-polar structure (Sandini and Metta, 2003; Boluda et al., 1996). Imaging system using Prism arrays  (Carles et al., 2016) or Risley-prism (Huang et al., 2021) expand the FoV by deflecting incoming rays and digitally super-resolving the overlapping central region, effectively producing foveated images through fusion. Similarly, a 3D-printed microlens system  (Thiele et al., 2017) combines pixels from four microlenses with different focal lengths to generate a foveated image. However, these methods hard-code the fovea region, necessitating mechanical camera motion to perform saccades.
More directly related to our work is a small body of literature on creating foveated imaging systems by locally correcting optical aberrations using dynamic wavefront shaping methods (Martinez et al., 2001). Such foveated images result from localized aberration correction, which can effectively address only one limited region of interest at a time. To correct off-axis aberrations locally, various optical solutions using different wavefront modulators have been proposed, including reflective SLMs (Wick et al., 2002; Curatu et al., 2005), transmissive SLM (Harriman et al., 2006), and reflective DMs (Zhao et al., 2008b, a). However, these methods are either bandwidth-limited and require polarized light (as with SLMs), or they require folded optical paths and complex system designs (as with reflective modulators), which is counterproductive for miniaturizing and simplifying optical systems. Furthermore, they typically optimize the optics for focusing at infinity and at a fixed oblique angle. In this work, we propose the use of a refractive wavefront modulator to enable compact and wide-band imaging. Additionally, we introduce a differential optics-based optimization method to tailor the desired wavefront across different imaging distances.
@@ 3. Wavefront Optimization with Differentiable Optics
In this section, we describe how to optimize DPP wavefront control patterns to (1) dynamically position the fovea anywhere within the image and (2) efficiently cover the field of view (FoV) with minimal saccades. We develop a forward model of the optical system shown in Fig. 1, comprising an achromatic doublet lens and a deformable phase plate, to enable ray tracing through the system. This model is implemented in a differentiable manner, allowing us to optimize the DPP’s wavefront parameters via back-propagation of the imaging loss. The subsections are organized as follows: Sec. 3.1 introduces the differentiable optical system for modeling and optimization; Sec. 3.2 details the DPP model within the system; Sec. 3.3 covers the optimization of a single DPP wavefront for foveated imaging at oblique angles; and Sec. 3.4 presents strategies for jointly optimizing multiple DPP wavefronts to cover the FoV for fovea stacking, under a limited number of images.
3.1. Differentiable Optics
Ray tracing through optical lenses allows for the calculation of point spread functions (PSFs) and supports realistic image simulation via ray tracing-based rendering. Recent advancements have made this process differentiable, enabling the optimization of optical lenses using ray-traced PSFs or simulated images as objectives (Wang et al., 2022; Yang et al., 2024). Built on the PyTorch auto-diff framework, DeepLens (Yang et al., 2024) performs ray-tracing by computing ray-surface intersections and refractions based on surface functions and their derivatives, and is fully differentiable for optimizing refractive lenses parameters. Unlike standard spherical or aspherical lenses, the DPP is a thin plate with freeform surface, requiring specialized modeling for differential optimization.
@@ 3.2. DPP Surface Modeling
where $\rho$ is normalized radial distance (normalized by aperture radius), $\varphi$ is the azimuthal angle, $Z_{k}$ and $w_{k}$ are the k-th Zernike polynomials and corresponding coefficients following OSA indices (Thibos et al., 2002). Partial derivatives with respect to $x$ and $y$ can be therefore derived and parameterized by $w_{k}$. More details on Zernike polynomials and their derivative calculation is provided in Supplement.
@@ 3.3. Optimized Wavefronts for Foveated Imaging
To maximize the imaging quality at a specific oblique angle, corresponding rays are sampled and traced through the optical system, consisting of a DPP and a lens. The PSF radius serves as the optimization loss, guiding the optimization of the DPP wavefront pattern via backpropagation. Specifically, for each sampled object point $P$ at distance $D$, $spp$ points $\{P_{i}|1\leq i\leq spp\}$ are sampled on the aperture plane to generate the sampling rays $\{R_{a}=\overset{\rightarrow}{PP_{i}}|1\leq i\leq spp\}$. Each sampled ray is then traced through the optical stack to reach the image plane at positions $p_{i\lambda}=RT(R_{i};\lambda)$, given a specific wavelength $\lambda$. The average distance of the traced image points from their center is considered the mean PSF radius, which serves as the loss function to be minimized via back-propagation. For color images, the PSF radius is further integrated over multiple wavelengths.
(2) $r(P)=\frac{1}{3}\sum_{\lambda}^{3}\frac{1}{spp}\sum_{i}^{spp}{|p_{i\lambda}-\bar{p_{\lambda}}|}$
Figure 2. Side view of the calibrated simulation system. The object plane is set 625 mm away. The imaging setup includes a DPP, a 50 mm achromatic doublet lens (Thorlabs AC-254-050A), and a Bayer pattern RGB sensor (FLIR GS3-U3-41C6C-C).
@@
Although aberration can be effectively corrected locally for different oblique angles $\phi$, the size of the fovea varies by location. In Fig 4, for each optimized DPP wavefront pattern, the MTF50 (defined as the sptail frequency at which the MTF drops to $50\%$) is visualized within a circular FoV with radius $R$, which circumscribes the square sensor (shown as a gray square). Given a specific sharpness threshold (e.g., 30 line pairs per mm), the radial and angular regions exceeding this threshold are identified and plotted as a rectangular region in a normalized radial-angular diagram. The normalized radius $\rho$ is related to the angle $\phi$ by: $\rho=f\tan(\phi)/R$, where $f$ is the effective focal length. The sharp coverage area tends to shrink as it approaches the periphery. It is estimated that approximately 100 images are needed to cover the entire square sensor region (shaded in gray in the radial-angular diagram) with resolution exceeding 30 lp/mm. More detailed analysis is provided in Supplementary. To capture sharp images across the full FoV more efficiently, a trade-off between sharpness and coverage must be considered. Notably, a secondary peak in sharpness often appears in the opposite direction (as shown at $\theta=\pm\pi$ in Fig. 4a), due to the known symmetries of certain aberrations, which can potentially be exploited.
Figure 5. Joint optimization of 5 DPP patterns using PSF grid stacking. (b) After optimization, each DPP deformation is specialized to optimize quality over different local regions while jointly covering the full FoV. (a) As an example the $1^{st}$ DPP pattern produces the best focus in the top-right and bottom-right regions. (c) The minimal value of PSF grid raidus $r_{min}$ across the stack serves as the optimization loss, while the mask index $n^{*}$ indicates which DPP pattern is best for any given region, which can be used to generate a composite PSF map for the image after stacking.
3.4. Optimize Wavefronts for Fovea Stacking

### v2

@@ Abstract.
The desire for cameras with smaller form factors has recently led to a push for exploring computational imaging systems with reduced optical complexity such as a smaller number of lens elements. Unfortunately such simplified optical systems usually suffer from severe aberrations, especially in off-axis regions, which can be difficult to correct purely in software.
In this paper we introduce Fovea Stacking, a new type of imaging system that utilizes an emerging dynamic optical component called the deformable phase plate (DPP) for localized aberration correction anywhere on the image sensor. By optimizing DPP deformations through a differentiable optical model, off-axis aberrations are corrected locally, producing a foveated image with enhanced sharpness at the fixation point - analogous to the eye’s fovea. Stacking multiple such foveated images, each with a different fixation point, yields a composite image free from aberrations. To efficiently cover the entire field of view, we propose joint optimization of DPP deformations under imaging budget constraints. Due to the DPP device’s non-linear behavior, we introduce a neural network-based control model for improved agreement between simulation and hardware performance.
@@
journal: TOGccs: Computing methodologies Computational photography
Figure 1. Fovea Stacking is a novel computational camera paradigm aimed at simplifying (and ultimately miniaturizing) camera optics. (a) The optical system consists of an achromatic doublet lens, which is highly aberrated, and a refractive deformable phase plate (DPP), which can be used to correct the aberrations for a local region of interest (the fovea). While this optical system cannot simultaneously correct all aberrations across the entire image, the DPP is capable of dynamically moving the fovea anywhere in the image. For example, the DPP deformation in (a) corrects aberrations along the red beam for the corresponding red box fovea on the sensor. (b) shows magnified crops of foveated images captured with different DPP deformations. The left column shows the image from (a), the other columns have the fovea moved to the other respective boxes. (c) By stacking up the sharpest regions from different foveated images (i.e. a fovea stack), an un-aberrated all-in-focus image can be recovered. Other applications such as object tracking and depth-dependent focusing are discussed in the text. :
@@ 1. Introduction
In this work we use the ability of the DPP to dynamically and locally shape an optical wavefront in a compact optical system to demonstrate a new type of foveated imaging system (see Fig. 1). The optical design consists of a combination of a highly aberrated achromatic double lens and a DPP that can be used for aberration correction. This optical system lacks the complexity to simultaneously correct for all aberrations everywhere in the image like a classical camera lens, however, it has the ability to correct for the aberrations in a localized region of interest (the fovea) and produce an image with excellent fidelity in that region. Moreover, through dynamic control of the DPP, this fovea region can be placed anywhere in the image plane using direct electrostatic actuation (c.f. Fig. 1b, Fig. 2). A stack of such foveated images, a.k.a. a Fovea Stack, can be used to reconstruct an unabberrated, high fidelity image (c.f. Fig. 1c).
@@
Figure 2. (a) physical layout of DPP device, composed of a deformable membrane, a optical liquid cavity, and a rigid glass substrate containing hexagonal electrodes. Figure adapted from (Rajaeipour et al., 2021). (b) localized aberration of different oblique angles can be corrected by dynamically changing DPP patterns. The green-channel wavefront aberration on exit-pupil plane and the corresponding colored spot diagrams are shown.
@@ 2.1. Wavefront Modulators
Recently, a novel optofluidic refractive wavefront modulator named a deformable phase plate (DPP) was proposed by (Banerjee et al., 2018). As shown in Fig. 2a, the DPP consists of a thin optically transparent, liquid-filled cavity sandwiched between a deformable membrane and a rigid glass surface. Its shape can be controlled by exerting electrostatic forces on the membrane via voltages applied to a hexagonal grid of 63 transparent electrodes on the rigid glass back panel. Using this principle, a DPP enables distortion correction up to seventh radial Zernike order (Rajaeipour et al., 2021). In this work, we adopt the DPP as a refractive wavefront modulator to create a compact imaging system. However, it is observed that traditional linear control strategies are inaccurate for large wavefront deviations. To address this, we propose a neural network-based control strategy to improve accuracy.
Compared to adaptive optics (AO) systems (Hampson et al., 2021; Wang et al., 2018), which typically correct for aberrations introduced by atmospheric turbulence or other external media before light reaches the imaging system, our method focuses on correcting distortions caused by the inherent imperfections of simple, compact imaging optics, tailored to the specific optical system in use.
@@ 2.2. Focus Stacking
Focus Stacking extends depth of field by capturing multiple images at varying focus distances. From this image stack, a sharp all-in-focus (AIF) image can be reconstructed. Similarly, Fovea Stacking aims to reconstruct a sharp, aberration corrected image from a stack of regionally sharp images. Focus Stacking can also be used correct for field curvature, where the focal region is a curved manifold instead of a plane. It can therefore serve as a baseline for our work.
Traditional multi-focus image fusion methods reconstruct all-in-focus images by identifying and compositing the sharpest regions from multiple input images  (Huang and Jing, 2007; Nayar and Nakagawa, 1994). This fusion can occur in the spatial domain or in transformed domains, such as Laplacian pyramids (Wang and Chang, 2011; Sun et al., 2018) and discrete wavelet transforms (Pradnya and Sachin, 2013). For comprehensive reviews of traditional and deep learning-based methods, see (Zafar et al., 2020; Zhang, 2021). Alternatively, some approaches jointly reconstruct both a depth map and an all-in-focus image - for example, Wang et al. (2021) use 3D convolution to learn attention weights from the focus stack, which are then used to produce both the AIF image and the corresponding depth map.
Traditional focus adjustment often involves mechanical movement of the sensor or lens elements, which can induce vibrations during large-amplitude, high-frequency motion (Kuthirummal et al., 2010). As an alternative, focus-tunable liquid lens (FTLs) have been adopted for focus-sweeping applications (Miau et al., 2013), but they are limited to tuning optical power without the ability to correct higher-order aberrations. Although the adjustable range of the DPP is limited, it is capable of adjusting the focus distance by modulating the ”defocus” term in the Zernike polynomial. Our method, therefore, supports stacking not only in the x-y direction but also along the depth direction, enabling aberration-corrected imaging for extended depth-of-field applications.
2.3. Differentiable Optics
Recent advances in computational optics demonstrate the effectiveness of end-to-end co-design of optics and reconstruction algorithms (Sitzmann et al., 2018; Sun et al., 2021). Central to this approach is a differentiable optics model. The first differentiable implementation of light propagation through a diffractive optical element (DOE) (Sitzmann et al., 2018) enabled applications in extended depth of field and super-resolution. This DOE model has since supported optical designs for wide field-of-view imaging (Peng et al., 2019), coded super-resolution SPAD imaging (Sun et al., 2020), and multi-aperture hyperspectral imaging (Shi et al., 2024). To support the end-to-end design of compound refractive lenses beyond simple wave-optics systems, researchers introduced both neural proxies (Tseng et al., 2021) and differentiable ray-tracing (Sun et al., 2021), which computes ray–surface intersections and refractions with full differentiability with respect to lens parameters. Subsequent work on differentiable ray-tracing reduced memory usage (Wang et al., 2022) and enabled automatic lens design from random initialization (Yang et al., 2024a), leading to the design of compound fluidic-freeform lenses (Na et al., 2024) and even hybrid refractive–diffractive lenses (Yang et al., 2024b).
In this work, we adopt a differentiable ray-tracing framework (Yang et al., 2024a) to optimize DPP deformation patterns for foveated imaging and stacking, targeting reconstruction quality. The customized differentiable model for the DPP device is detailed in Sec. 3.
@@ 2.4. Foveated Imaging
Foveated imaging, inspired by human vision, concentrates high resolution where it is most needed - similar to the function of the fovea. Many different approaches have been proposed to achieve spatially varying resolution through specialized sensor design or image fusion techniques. Retina-like sensors, for example, arrange photoreceptors in a spatially varying log-polar structure (Sandini and Metta, 2003; Boluda et al., 1996). Imaging systems using prism arrays  (Carles et al., 2016) or a Risley-prism (Huang et al., 2021) expand the FoV by deflecting incoming rays and digitally super-resolving the overlapping central region, effectively producing foveated images through fusion. Similarly, a 3D-printed microlens system  (Thiele et al., 2017) combines pixels from four microlenses with different focal lengths to generate a foveated image. However, these methods hard-code the fovea region, necessitating mechanical camera motion to perform saccades.
More directly related to our work is a small body of literature on creating foveated imaging systems by locally correcting optical aberrations using dynamic wavefront shaping methods (Martinez et al., 2001). Such foveated images result from localized aberration correction, which can effectively address only one limited region of interest at a time. To correct off-axis aberrations locally, various optical solutions using different wavefront modulators have been proposed, including reflective SLMs (Wick et al., 2002; Curatu et al., 2005), transmissive SLMs (Harriman et al., 2006), and reflective DMs (Zhao et al., 2008b, a). However, these methods are either bandwidth-limited and require polarized light (as with SLMs), or they require folded optical paths and complex system designs (as with reflective modulators), which is counterproductive for miniaturizing and simplifying optical systems. Furthermore, they typically optimize the optics for focusing at infinity and at a fixed oblique angle. In this work, we propose the use of a refractive wavefront modulator to enable compact and wide-band imaging. Additionally, we introduce a differential optics-based optimization method to tailor the desired wavefront across different imaging distances.
@@ 3. Wavefront Optimization with Differentiable Optics
In this section, we describe how to optimize DPP wavefront control patterns to (1) dynamically position the fovea anywhere within the image and (2) efficiently cover the field of view (FoV) with minimal saccades. We develop a forward model of the optical system shown in Fig. 1, comprising an achromatic doublet lens and a deformable phase plate, to enable ray tracing through the system. This model is implemented in a differentiable manner, allowing us to optimize the DPP’s wavefront parameters via back-propagation of the imaging loss. The subsections are organized as follows: Sec. 3.1 details the DPP model within differentiable optics system; Sec. 3.2 covers the optimization of a single DPP wavefront for foveated imaging at oblique angles. Sec. 3.3 presents strategies for jointly optimizing multiple DPP wavefronts to cover the FoV for Fovea Stacking, under a limited number of images.
@@ 3.1. DPP Surface Modeling
where $\rho$ is normalized radial distance (normalized by aperture radius), $\varphi$ is the azimuthal angle, $Z_{k}$ and $w_{k}$ are the k-th Zernike polynomials and corresponding coefficients following OSA indices (Thibos et al., 2002). Partial derivatives with respect to $x$ and $y$ can be therefore derived and parameterized by $w_{k}$. Because horizontal and vertical tilt only cause geometric distortion – resulting in unwanted pixel misalignment across images rather than improved image sharpness – they are excluded from Zernike coefficient optimization and fixed at zero. Notice that we do not correct for radial lens distortion since it remains consistent across a stack of images as long as the DPP does not introduce tilt. More details on Zernike polynomials and their derivative calculation is provided in the Supplement.
@@ 3.2. Optimized Wavefronts for Foveated Imaging
To maximize the imaging quality at a specific oblique angle, corresponding rays are sampled and traced through the optical system, consisting of a DPP and a lens. The RMS spot size serves as the optimization loss, guiding the optimization of the DPP wavefront pattern via backpropagation. Specifically, for each sampled object point $P$ at distance $D$, $N_{a}$ points $\{P_{i}|1\leq i\leq N_{a}\}$ are sampled on the aperture plane to generate the sampling rays $\{R_{i}=Ray(P\rightarrow P_{i})|1\leq i\leq N_{a}\}$. Each sampled ray is then traced through the optical stack to reach the image plane at positions $p_{i\lambda}=RT(R_{i};\lambda)$, given a specific wavelength $\lambda$. The root-mean-square radial deviation of traced rays from their center, termed the RMS spot size, is used as the loss function minimized via backpropagation. As RMS spot size is very closely related to the area under the modulation transfer function MTF (Luc, 2022), it is a popular loss in lens design. For color images, the RMS spot size is further integrated over multiple wavelengths.
[…]
Figure 3. Side view of the calibrated simulation system. The object plane is set 652 mm away. The imaging setup includes a DPP, a 50 mm achromatic doublet lens (Thorlabs AC-254-050A), and a Bayer pattern RGB sensor (FLIR GS3-U3-41C6C-C).
@@
Although aberration can be effectively corrected locally for different oblique angles $\phi$, the size of the fovea varies by location. In Fig 5, for each optimized DPP wavefront pattern, the MTF50 (defined as the spatial frequency at which the MTF drops to $50\%$) is visualized within a circular FoV with radius $R$, which circumscribes the square sensor (shown as a gray square). Given a specific sharpness threshold (e.g., 30 line pairs per mm), the radial and angular regions exceeding this threshold are identified and plotted as a rectangular region in a normalized radial-angular diagram. The normalized radius $\rho$ is related to the angle $\phi$ by: $\rho=f\tan(\phi)/R$, where $f$ is the effective focal length. The sharp coverage area tends to shrink as it approaches the periphery. It is estimated that approximately 100 images are needed to cover the entire square sensor region (shaded in gray in the radial-angular diagram) with resolution exceeding 30 lp/mm. More detailed analysis is provided in Supplementary. To capture sharp images across the full FoV more efficiently, a trade-off between sharpness and coverage must be considered. Notably, a secondary peak in sharpness often appears in the opposite direction (as shown at $\theta=\pm\pi$ in Fig. 5a), due to the known symmetries of certain aberrations, which can potentially be exploited.
Figure 6. Joint optimization of 5 DPP patterns using grid stacking. (b) After optimization, each DPP deformation is specialized to optimize quality over different local regions while jointly covering the full FoV. (a) As an example the $1^{st}$ DPP pattern produces the best focus in the top-left and bottom-right regions. (c) The minimal value of RMS spot size $r_{min}$ across the stack serves as the optimization loss, while the mask index $n^{*}$ indicates which DPP pattern is best for any given region, which can be used to generate (d) a composite PSF map for the image after stacking.
3.3. Optimized Wavefronts for Fovea Stacking